# TE-NeXt: A LiDAR-Based 3D Sparse Convolutional Network for Traversability Estimation

### v2

@@ IV-D3 Performance in natural environments
In addition to a high mean intersection over union (mIoU) for traversability and non-traversability, the model achieves a true positive rate (TPR) of 81.3%, indicating a robust capacity to accurately identify traversability areas. Concurrently, it exhibits a high true negative rate (TNR) of 91.8%, which validates its efficacy in excluding non-traversability regions. In most of the metrics obtained, the TE-NeXt model demonstrates superior performance, although it is sometimes surpassed by the methods evaluated in two-dimensional space, which, as previously stated, exhibit significantly diminished performance when assessed through the same 3D criteria as the other methods.
In regard to methods dedicated to traversability estimation, such as nn Risk [60] and ScaTE [22], which is at the state of the art as evidenced in Table I, TE-NeXt is superior in the single available metric, as the code is not accessible.
Figure 5 presents the precision-recall curves for the evaluated traversability estimation methods, providing a comprehensive visualization of their performance across different classification thresholds. This graphical representation offers valuable insights into the trade-offs between precision and recall for each method, particularly in identifying non-traversable areas. As observed in the figure, the proposed TE-NeXt method (blue line) exhibits superior performance. It maintains high precision across a wide range of recall values, suggesting robust identification of non-traversable areas with minimal false positives. This characteristic is crucial for traversability estimation in autonomous navigation, where maintaining high precision (minimizing false positives) is often prioritized to ensure safety. The area under the precision-recall curve (AUC-PR), provides a quantitative measure for the overall performance comparison. Based on the visual inspection of the curves, TE-NeXt presents a high edge in this regard, suggesting its potential as a robust solution for traversability estimation in autonomous navigation systems.
TABLE VI: Traversability estimation performance of semantic segmentation methods in natural environments.
@@
SalsaNext [52] $\checkmark$ 73.5 90.0 73.3 68.9 95.5
SalsaNext [52] $\checkmark$ 70.9 85.0 70.5 65.4 95.0
@@ IV-E Qualitative Evaluation
Figure 6 presents the results in a visual form, where the different colours refer to the success or failure cases for each class considered: true positives (blue), true negatives (red), false positives (purple), false negatives (yellow). The complete inference of the test sequences belonging to the aforementioned datasets is publicly available on YouTube [^1] https://www.youtube.com/watch?v=mn6NKlnxjOs&t=122s.
Figure 6a illustrates the predictions of the proposed architecture in a cloud in an unstructured environment. It can be observed that the points classified by the network as traversable but whose label is not (false positives) are located at the limits of the different geometries detected by the LiDAR. This error is attributed to the very nature of the algorithm itself, which discretises the environment into voxels. With regard to the false negatives, two scenarios emerge: firstly, those that present evidence in the right section of the image of the diffuse geometry of the environment. Secondly, the false negatives that appear in the left section of the image are indicative of an error in the labelling of the data, given that the traversability of this part of the environment is questionable.
Figure 6b shows the predictions of the proposed architecture on a cloud in an urban environment. In this type of environment, as can be seen in the interpretation of results, the inference of the network is much more accurate, there are hardly any errors, and sources of error such as false positives occur for the same reasons mentioned above.
IV-F Fully Autonomous Navigation Framework
@@ IV-G Implementation details
Table VII provides an informative comparison of key efficiency metrics for the methods compared above. The results highlight the substantial differences in model size, memory needs and latency that can arise from architecture choices.With regard to the model size, it is evident that TE-NeXt fulfils one of the primary objectives of its design: the reduction of the model’s complexity without compromising its performance.
@@
V Conclusion
Future work will include the addition of visual information through a LiDAR-camera calibration, which will allow visual features to be introduced into the sparse tensors. Possibly with the aim of mitigating the main problem of supervised learning, the content of the labelled data, we will investigate the exposure of a hybrid method that allows continuous estimation of traversability.
TABLE VII: Comparison of semantic segmentation methods in terms of efficiency
@@ V Conclusion
VI Acknowledgments

### v3

@@ 5.2 Performance in natural environments
SalsaNext [51] $\checkmark$ 73.5 88.0 73.3 68.9 95.5
SalsaNext [51] $\checkmark$ 70.9 85.3 70.5 65.4 95.0
@@
Table 5: Traversability estimation performance of semantic segmentation methods in natural environments.
In addition to a high mean intersection over union (mIoU) for traversability and non-traversability, the model achieves a true positive rate (TPR) of 81.3%, indicating a robust capacity to accurately identify traversability areas. Concurrently, it exhibits a high true negative rate (TNR) of 91.8%, which validates its efficacy in excluding non-traversability regions. In most of the metrics obtained, the TE-NeXt model exhibits superior performance, however, it is occasionally outperformed by methods evaluated in two-dimensional space, which demonstrate a considerable degree of variation in results between different metrics, such as TPR and TNR. This suggests that these methods are not consistent for this specific type of environment. Additionally, as previously stated, they demonstrate significantly lower performance when evaluated using the same 3D criteria as the other methods.
In regard to methods dedicated to traversability estimation, such as nn Risk [59] and ScaTE [22], which is at the state of the art as evidenced in Table 1, TE-NeXt is superior in the single available metric, as the code is not accessible.
As the focus of this paper is on performance in unstructured environments, Figure 4 presents the precision-recall curves for the evaluated traversability estimation methods, providing a comprehensive visualization of their performance across different classification thresholds. This graphical representation offers valuable insights into the trade-offs between precision and recall for each method, particularly in identifying non-traversable areas. As observed in the figure, the proposed TE-NeXt method (blue line) exhibits superior performance. It maintains high precision across a wide range of recall values, suggesting robust identification of non-traversable areas with minimal false positives. This characteristic is crucial for traversability estimation in autonomous navigation, where maintaining high precision (minimizing false positives) is often prioritized to ensure safety. The area under the precision-recall curve (AUC-PR), provides a quantitative measure for the overall performance comparison. Based on the visual inspection of the curves, TE-NeXt presents a high edge in this regard, suggesting its potential as a robust solution for traversability estimation in autonomous navigation systems.
@@ 5.3 Qualitative Evaluation
Figure 5(a) illustrates the outcomes of the proposed architecture in an unstructured environment. It can be observed that the points classified by the network as traversable but whose label is not (false positives) are located at the limits of the different geometries detected by the LiDAR. This error is attributed to the very nature of the algorithm itself, which discretises the environment into voxels. With regard to the false negatives, two scenarios emerge: firstly, those that present evidence in the right section of the image of the diffuse geometry of the environment. Secondly, the false negatives that appear in the left section of the image are indicative of an error in the labelling of the data, given that the traversability of this part of the environment is questionable.
Figure 5(b) shows the predictions of the proposed architecture in an urban environment. In this type of environment, as can be seen in the interpretation of results, the inference of the network is much more accurate, there are hardly any errors, and sources of error such as false positives occur for the same reasons mentioned above.
5.4 Fully Autonomous Navigation Framework
@@ 5.5 Implementation details
Table 6 provides an informative comparison of key efficiency metrics for the methods compared above. The results highlight the substantial differences in model size, memory needs and latency that can arise from architecture choices. With regard to the model size, it is evident that TE-NeXt fulfils one of the primary objectives of its design: the reduction of the model’s complexity without compromising its performance.
@@
Table 6: Comparison of semantic segmentation methods in terms of efficiency.
6 Conclusion
Future work will include the addition of visual information through a LiDAR-camera calibration, which will allow visual features to be introduced into the sparse tensors. Possibly with the aim of mitigating the main problem of supervised learning, the annotation process, we will investigate the exposure of a hybrid method that allows continuous estimation of traversability.
7 Acknowledgments